# End-to-End Demonstration for CubeSatellite Quantum Key Distribution

### v1

@@ 3 System and Result
Figure 1: End to End Satellite Free Space QKD Testbench. (a) The test platform is divided into three parts, with red parallel lines indicating the quantum signal at a wavelength of 785 nm, purple indicating the beacon optical channel at a wavelength of 905 nm, and green dashed lines indicating the distribution of the electrical signal. The middle is the communication channel. The right side is the receiver deployed on the ground. (b) is the field programmable gate arrays (FPGA) board and the adapted quantum source which is coupled by single mode fiber. (c) is the beacon driver which is triggered by FPGA, the laser diode is on the right edge of the board. The modulated pulses have a width of less than 4.1ns and a repetition of 100 kHz to achieve the goal of high SNR and low average power. (d) is the receiving box in Bob side, the top left 3 waveplates are used for polarisation compensation while the 4th is the bandpass filter to compress the noise. The optical path is marked using red line in the figure. (e) is the SPCMs used for single photon detection, the bottom is the time tagger developed by Bristol which has 16 parallel channels with 1.6 ps resolution [32].
@@
This paper demonstrates a free-space QKD validation platform consisting of a transmitter (Alice) based on the BB84 protocol, a transmission channel with controlled background noise and loss, a receiver (Bob), and a classical communication link for completing post-processing. The platform is used to validate the functional operation of the QKD system in the laboratory, the performance at different losses and noise, and the post-processing algorithms. The transmitter used in the platform is the prototype for the ROKs mission (led by Craft Prospect Ltd) and a UK-Canada mission (piggy backed on the QEYSSAT mission). An upgraded version is being adapted for the UK EPSRC Quantum Communication Hub mission [18], which is scheduled to launch into space for in-orbit demonstration in 2024. The receivers will be deployed on optical ground stations behind suitable telescopes depending on the mission. Fig. 1 shows the testbench diagram and specific design for the source and receiver. In (a) the left side is the transmitter, which consists of PC, FPGA, a quantum source, a beacon source and the corresponding optical path. The ZYNQ7000 FPGA provides a pseudorandom binary sequence (PRBS) generator for triggering the quantum source emitting the different polarisation states in a pseudo random sequence. The QKD source uses 4 individual laser diodes to emit pulses with different polarisation. A 2 nm bandpass filter is used to minimise the spectrum differences between different diodes to avoid potential security loopholes. The driver current is configured separately for each laser diode to make sure the intensity and pulse width is identical with a reasonable difference, as shown in Table. 2. An HDBC encoder is implemented to trigger the beacon source generating a 100 kHz bright laser signal, which is to achieve dual-channel synchronisation with jitter of less than 12 ps. A variable optical attenuator is connected to the quantum source via single mode fiber to set the $\mu$ which then is collimated into free space. The beacon source is designed based on a multimode laser diode with ultra high peak power up to 100 W. An aspheric lens is used for preliminary collimation and the ND filter is used for simulating the internal loss by aperture clipping due to the multimode and anisotropic beam. Finally a 820 nm cutoff dichroic mirror is used for combining the two beams (beacon and quantum) into a co-propagating beam. In the transmission channel, a variable ND filter is placed on the propagation path of the beacon to simulate the loss variations during the satellite’s orbit. The background light in the system comes from the illumination of the sunlight and the intensity is controlled by a curtain. It should be noted that this is the prototype system demonstration before the development of flight model for final launch. We use it here mainly to prove operation at in high loss scenarios and subsequent secure key generation. Our ongoing device level work is proving that SWAP for the cubesatellite and other space environment requirements can be met [33].
@@
In the receiving block located on the right, another 820 nm dichroic mirror is used to separate the quantum and beacon beams. The quantum signal is split into four channels by the receiver box and coupled to the corresponding SPCMs via a multimode fibre. A three waveplate array at the box entrance is used to correct for polarisation errors between the satellite and the OGS, and a bandpass filter with a transmission efficiency of 50% suppresses all spectral light except the 785 nm with a 10 nm window to improve the SNR. Inside the receiving box, a beamsplitter (BS) splits the quantum signal into the two measurement bases, one arm of which passes through another waveplate used to rotate to the diagonal basis so that an H-V polarising BS can discriminate D-A. For the beacon detection, a detector circuit developed based on the Hamamatsu avalanche photodiode (APD) S13282 is deployed with a sensitivity of 4 MV/W and a noise level of 50 $fW/\sqrt{Hz}$ so that the faint pulse after up to 80 dB loss still can be detected. The two computers used to control Alice and Bob respectively are connected via TCP/IP protocol via a network switch for the necessary classical communication and post-processing to generate security keys. Table. 2 shows the parameters set in the system, which are as consistent as possible with the parameters used in future practical tasks to achieve more accurate results.
@@ 3.1 Channel Loss Capability
Figure 2: SKR/QBER Vs channel loss. The system loss is 9.6 dB which includes all the components both in transmitter (1 dB) and receiver (8.6 dB). The additional channel loss is configured from 6.3 dB to 37.3 dB using a variable optical ND filter. Adding the system loss and the channel loss takes us to a maximum of 46.9 dB, which is beyond the maximum loss estimated in Section. 2. The figure shows the SKR change with the channel loss increase with the parameters configured as table. 2. The initial trend before 30 dB is linear and then decreasing faster as the QBER increases rapidly. [34]
@@ 3.2 Background Noise Effect
Figure 3: SKR / QBER Vs background noise. The background brightness is characterized using the noise count rate of the SPCMs. The hollow legends represent SKR under different background light noise, and the solid legends represent QBER (shaded ares are indicative of the scale of error in parameters). The system performance using three different values of additional channel loss (10 dB, 16 dB, and 25 dB) is experimentally studied. System parameters are configured in table. 2
The QBER and SKR under different background noise are investigated to see how the environmental light intensity will affect the performance, which is characterized by the background count rate detected using SPCMs. In Fig. 3, under the three different additional channel loss, SKR and QBER remain essentially unchanged initially. However, as the noise level further increases, QBER rises sharply after a certain point, accompanied by a sharp drop in SKR. Increased channel loss means that the noise level remains unchanged but the signal intensity is attenuated while here the noise level increases and the signal intensity remains unchanged. It is worth noting that the inflection point of SKR with different losses occurs at different background noise levels, 30 kbps, 200 kbps, 1000 kbps respectively. But this result is consistent with the trend of QBER and SKR when the loss increases in the Fig. 1 if considering from the point of view of SNR discussed in the section 3.4.
@@ 3.3 Effect of Gate Width
Figure 4: Gate width effect on SKR. The left figure shows the SKR and QBER change with channel loss using different gate widths to do postprocessing. The solid legends correspond to the SKR (left axis) while the hollow legends represents the QBER (right axis). The right figure shows the performance change under different gate widths. The hollow legends presents the normalised SKR (NSKR) while the solid legends present the QBER. The parameters used are configured in table. 2
@@ 3.4 QBER and ESNR
Figure 5: QBER Vs ESNR. The ESNR combine the background noise, $\mu$, channel loss and gate width together, which evaluates the performance of QKD more efficiently. The right figure shows the QBER from different cases (Different channel losses, background noise, and $\mu$) shows the same trends when increasing the ESNR.
From the Fig. 5, the left figure shows the composition of the photon events registered by SPCMs, the respective contributions of signal and noise photons to them and the role of the gating window. When a narrow window is applied, less noise is introduced but more signal is lost and vice versa. The graph on the right depicts how the QBER grows for different system settings with variations in ESNR. In the first five curves, The background noise level and the channel losses are the same but the $\mu$ are set to 0.1, 0.3, 0.5, 0.8, 1.0 respectivelyIt should be noted that the $\mu$ here is not set based on the theoretical optimised value for maximising the security key rate. Here we mainly want to show that the ESNR is more reliable for qualifying the performance so that the $\mu$ value is picked uniformly in the range of 0.1 - 1.0 from the engineering point of view.. the 6 - 8 curves show the experimental result with different background noise, different channel losses (10 dB, 16 dB, 25 dB) and same $\mu=0.1$. With the equation 6, the QBER keep the same falling trend with the ESNR increasing. This gives an interesting point that when analysing the effect on QBER from different losses, $\mu$, background noise, gating widths, each parameter produces a different trend. However, when we transform all the above parameters into ESNR and look at it’s relationship with QBER, all these points from different data sets (with different parameters) obey the same variation. This means that using ESNR to estimate and evaluate QBER and SKR can lead to more valid conclusions.
@@ 4 Improvement and Optimisation
Loss budget. In the quantum channel of the end-to-end testbench, the finite optical path length introduces no geometric loss to the system, the 7.4 dB attenuation is caused by the fiber coupling efficiency and the bandpass filter’s transmission, called $\eta_{I}$. The narrowband filter has a transmission of 50% for 785 nm quantum signal, contributing 3 dB loss. The remainder comes mainly from free space-to-multimode fibre coupling at the output of the receiver box and a pair of dichroic mirrors with a transmission of approximately 80%. The maximum channel loss added in the experiment is 37.3 dB, considering the 2.2 dB (60%) detection efficiency of the SPCMs, the total loss that can be tolerated is 46.9 dB in this work. In the real scenario, the maximum geometric channel loss is about 27.7 dB with a 10 microrad divergence and 70 cm telescope aperture. With 7.9 dB absorption loss, the total loss budget with the current setup is $Loss_{Q}=45.8dB$ (Hub mission). The improvement of bandpass filter transmission in the OGS from 50% to 90% will reduce the aggregated loss by 2.6 dB. For the beacon and synchronisation channel, the maximum loss budget is 79.4 dB for the ROKs cube satellite mission due to a larger divergence (100 microrad) as it is also used for the pointing and tracking, while it is only 50.0 dB for the hub mission (synchronisation only). In this test, the beacon not only passes through the same ND filter with the quantum beam, but also through an extra 42.2 dB ND filter (calibrated at the wavelength of transmission). The dichroic mirror pair and focusing lens in front of the APD contribute another 3 dB. So the total loss suffered in this experiment is 85.2 dB, which clearly suggests it can cope with operation in orbit for the ROKs mission. The Hubs mission will utilise a single mode laser which can provide a 400 mW peak power but with a small divergence of 18.7 microrad so that the received intensity could provide enough SNR ($>10$) for timing and synchronisation.
@@
Dark count. The DCR of the detector used here is 500 Hz, and using SPCMs with DCR $<$ 100 Hz will result in a 7 dB increase in SNR. The reduced DCR will further decrease the QBER and increase the secuirity key rate.
Finite key length. LEO’s intermittent communication window allows a limited length of key to be generated per satellite pass, which exacerbates the generation of secure keys as a significant portion of received signal must be sacrificed for accurate parameter estimation while retaining enough raw key for there to be a finite secure key after postprocessing. Considering the $QBER_{I}$ = 0.5%, the finite key length is about 6.5 Mbits with the elevation angle larger than 10 degrees. [34]. The aimed $R_{Q}$ is 400 MHz which could get at least a 12 dB SKL improvement over this 25 MHz system (assuming comparable $QBER$). Improving the quantum efficiency of the detector is another way to improve SKL but wouldn’t be too much.
@@ Appendix A Channel Loss Estimation
Figure 6: The figure shows the distance and elevation angle change during the visibility window. The circle line shows the Elevation angle while the blue solid line shows the distance. The result is modeling based on the 3D Doppler model for the LEO [21]. The result is calculated with the value in the table LABEL:tab3.
Here, the channel loss modeling is given based on the diffraction limit. According to the analysis of the visibility window and the link efficiency, the elevation angle is configured larger than 10 degrees. The corresponding minimal and maximal distance is 500 km and 1700 km. Based on the profile of the Gaussian beam, the diffraction limited half angle with the threshold of $1/e^{2}$ is given by

### v2

@@ 3 System and Experimental Result
Figure 1: End to End Satellite Free Space QKD Testbench. (a) The test platform is divided into three parts, with red parallel lines indicating the quantum signal at a wavelength of 785 nm, purple indicating the beacon optical channel at a wavelength of 905 nm, and green dashed lines indicating the distribution of the electrical signal. The middle is the communication channel. The right side is the receiver deployed on the ground. (b) is the field programmable gate arrays (FPGA) board and the adapted quantum source which is coupled by single mode fiber. (c) is the beacon driver which is triggered by FPGA, the laser diode is on the right edge of the board. The modulated pulses have a width of less than 4.1ns and a repetition of 100 kHz to achieve the goal of high SNR and low average power. (d) is the receiving box in Bob side, the top left 3 waveplates are used for polarisation compensation while the 4th is the bandpass filter to compress the noise. The optical path is marked using red line in the figure. (e) is the SPCMs (Excelitas SPCM-NIR) used for single photon detection, the bottom is the time tagger developed by Bristol which has 16 parallel channels with 1.6 ps resolution [33].
@@
This paper demonstrates a free-space QKD validation platform consisting of a transmitter (Alice) based on the BB84 protocol, a transmission channel with controlled background noise and loss, a receiver (Bob), and a classical communication link for completing post-processing. The platform is used to validate the functional operation of the QKD system in the laboratory, the performance at different losses and noise, and the post-processing algorithms. The transmitter used in the platform is the prototype for the ROKs mission (led by Craft Prospect Ltd) and a UK-Canada mission (piggy backed on the QEYSSAT mission). An upgraded version is being adapted for the UK EPSRC Quantum Communication Hub mission [19], which is scheduled to launch into space for in-orbit demonstration in 2024. The receivers will be deployed on optical ground stations behind suitable telescopes depending on the mission. Fig. 1 shows the testbench diagram and specific design for the source and receiver. In (a) the left side is the transmitter, which consists of PC, FPGA, a quantum source, a beacon source and the corresponding optical path. The ZYNQ7000 FPGA provides a pseudorandom binary sequence (PRBS) generator for triggering the quantum source emitting the different polarisation states in a pseudo random sequence. The QKD source uses 4 individual laser diodes to emit pulses with different polarisation. A 2 nm bandpass filter is used to minimise the spectrum differences between different diodes to avoid potential security loopholes. The driver current is configured separately for each laser diode to make sure the intensity and pulse width is identical with a reasonable difference, as shown in Table. 1. An HDBC encoder is implemented to trigger the beacon source generating a 100 kHz bright laser signal, which is to achieve dual-channel synchronisation with jitter of less than 12 ps. A variable optical attenuator is connected to the quantum source via single mode fiber to set the $\mu$ which then is collimated into free space. The beacon source is designed based on a multimode laser diode with ultra high peak power up to 100 W. An aspheric lens is used for preliminary collimation and the ND filter is used for simulating the internal loss by aperture clipping due to the multimode and anisotropic beam. Finally a 820 nm cutoff dichroic mirror is used for combining the two beams (beacon and quantum) into a co-propagating beam. In the transmission channel, a variable ND filter is placed on the propagation path of the beacon to simulate the loss variations during the satellite’s orbit. The background light in the system comes from the illumination of the sunlight and the intensity is controlled by a curtain. It should be noted that this is the prototype system demonstration before the development of flight model for final launch. We use it here mainly to prove operation at in high loss scenarios and subsequent secure key generation. Our ongoing device level work is proving that SWAP for the cubesatellite and other space environment requirements can be met [34]. The QKD payload shown in Fig. 1 (b) has three layers, including FPGA core board, source driving board and power board. Each layer has a area about 10cm x 10cm equivalently and the next integrated version has proofed that the whole payload can be fit into a 0.5 U cage.
@@
In the receiving block located on the right, another 820 nm dichroic mirror is used to separate the quantum and beacon beams. The quantum signal is split into four channels by the receiver box and coupled to the corresponding SPCMs via a multimode fibre. A three waveplate array at the box entrance is used to correct for polarisation errors between the satellite and the OGS, and a bandpass filter with a transmission efficiency of 50% suppresses all spectral light except the 785 nm with a 10 nm window to improve the SNR. Inside the receiving box, a beamsplitter (BS) splits the quantum signal into the two measurement bases, one arm of which passes through another waveplate used to rotate to the diagonal basis so that an H-V polarising BS can discriminate D-A. For the beacon detection, a detector circuit developed based on the Hamamatsu avalanche photodiode (APD) S13282 is deployed with a sensitivity of 4 MV/W and a noise level of 50 $fW/\sqrt{Hz}$ so that the faint pulse after up to 80 dB loss still can be detected. The two computers used to control Alice and Bob respectively are connected via TCP/IP protocol via a network switch for the necessary classical communication and post-processing to generate security keys. Table. 2 shows the parameters set in the system, which has some difference with the actual mission plan, but the result could be convert to the mission configuration as the estimation of proof-principle.
@@ 3.1 Channel Loss Capability
Figure 2: Experimental result of SKR/QBER Vs channel loss. The system loss is 7.4 dB which includes all the components both in transmitter  [^5] In the real scenario, the mean photon number is set at the output aperture of the satellite telescope, so there is no extra loss introduced. (1 dB) and receiver (6.4 dB). The additional channel loss is configured from 6.3 dB to 37.7 dB using a variable optical ND filter. Adding the system loss and the channel loss takes us to a maximum of 47.3 dB, which is beyond the maximum loss estimated in Section. 2. The figure shows the SKR change with the channel loss increase with the parameters configured as table. 2. The initial trend before 30 dB is linear and then decreasing faster as the QBER increases rapidly [35]. The background noise is 2.7 kHz in the experiment.
@@ 3.2 Background Noise Effect
Figure 3: Experimental result of SKR / QBER Vs background noise. The background brightness is characterized using the noise count rate of the SPCMs when sending no quantum light. The hollow legends represent SKR under different background light noise, and the solid legends represent QBER (shaded ares are indicative of the scale of error in parameters). The system performance using three different values of additional channel loss (10 dB, 16 dB, and 25 dB) is experimentally studied. System parameters are configured in table. 1
The QBER and SKR under different background noise are investigated to see how the environmental light intensity will affect the performance. The background noise count rate is measured using the receiver unit while sending no quantum light. The background noise intensity is controlled by adjusting the light leakage into the (covered) experimental environment from external room/day light. In Fig. 3, under the three different additional channel loss, SKR and QBER remain essentially unchanged initially. However, as the noise level further increases, QBER rises sharply after a certain point, accompanied by a sharp drop in SKR. Increased channel loss means that the noise level remains unchanged but the signal intensity is attenuated while here the noise level increases and the signal intensity remains unchanged. It is worth noting that the inflection point of SKR with different losses occurs at different background noise levels, 30 kbps, 200 kbps, 1000 kbps respectively. But this result is consistent with the trend of QBER and SKR when the loss increases in the Fig. 1 if considering from the point of view of SNR discussed in the section 3.4.
@@ 3.3 Effect of Gate Width
Figure 4: Experimental result of gate width effect on SKR. The left figure shows the SKR and QBER change with channel loss using different gate widths to do postprocessing. The solid legends correspond to the SKR (left axis) while the hollow legends represents the QBER (right axis). The right figure shows the performance change under different gate widths. The hollow legends presents the normalised SKR (NSKR) while the solid legends present the QBER. The parameters used are configured in table. 2 and the extra background noise is 400 kHz except the DCR.
@@ 3.4 QBER and ESNR
Figure 5: Experimental result of QBER Vs ESNR. The ESNR combine the background noise (BK), $\mu$, channel loss and gate width together, which evaluates the performance of QKD more practically. The right figure shows the QBER from different cases (Different channel losses, background noise, and $\mu$) shows the same trends when increasing the ESNR.
From the Fig. 5, the left figure shows the composition of the photon events registered by SPCMs, the respective contributions of signal and noise photons to them and the role of the gating window. When a narrow window is applied, less noise is introduced but more signal is lost and vice versa. The graph on the right depicts how the QBER grows for different system settings with variations in ESNR. In the first five curves, The background noise level and the channel losses are the same but the $\mu$ are set to 0.1, 0.3, 0.5, 0.8, 1.0 respectively [^6] It should be noted that the $\mu$ here is not set based on the theoretical optimised value for maximising the security key rate. Here we mainly want to show that the ESNR is more reliable for qualifying the performance so that the $\mu$ value is picked uniformly in the range of 0.1 - 1.0 from the engineering point of view.. the 6 - 8 curves show the experimental result with different background noise, different channel losses (10 dB, 16 dB, 25 dB) and same $\mu=0.1$. With the equation 6, the QBER keep the same falling trend with the ESNR increasing. This gives an interesting point that when analysing the effect on QBER from different losses, $\mu$, background noise, gating widths, each parameter produces a different trend. However, when we transform all the above parameters into ESNR and look at it’s relationship with QBER, all these points from different data sets (with different parameters) obey the same variation. This means that using ESNR to estimate and evaluate QBER and SKR can lead to more valid conclusions.
@@ 4 Improvement and Optimisation
Loss budget. In the quantum channel of the end-to-end testbench, the finite optical path length introduces no geometric loss to the system, the 7.4 dB attenuation is caused by the fiber coupling efficiency and the bandpass filter’s transmission, called $\eta_{I}$. The narrowband filter has a transmission of 50% for 785 nm quantum signal, contributing 3 dB loss. The remainder comes mainly from free space-to-multimode fibre coupling at the output of the receiver box and a pair of dichroic mirrors with a transmission of approximately 80%. The maximum channel loss added in the experiment is 37.3 dB, considering the 2.2 dB (60%) detection efficiency of the SPCMs, the total loss that can be tolerated is 46.9 dB in this work. In the real scenario, the maximum geometric channel loss is about 27.7 dB with a 10 microrad divergence and 70 cm telescope aperture. With 7.9 dB absorption loss, the total loss budget with the current setup is $Loss_{Q}=45.8dB$ (SPOQC Mission). The improvement of bandpass filter transmission in the OGS from 50% to 90% will reduce the aggregated loss by 2.6 dB. For the beacon and synchronisation channel, the maximum loss budget is 79.4 dB for the ROKs cube satellite mission due to a larger divergence (100 microrad) as it is also used for the pointing and tracking, while it is only 50.0 dB for the hub mission (synchronisation only). In this test, the beacon not only passes through the same ND filter with the quantum beam, but also through an extra 42.2 dB ND filter (calibrated at the wavelength of transmission). The dichroic mirror pair and focusing lens in front of the APD contribute another 3 dB. So the total loss suffered in this experiment is 85.2 dB, which clearly suggests it can cope with operation in orbit for the ROKs mission. The SPOQC mission will utilise a single mode laser which can provide a 400 mW peak power but with a small divergence of 18.7 microrad so that the received intensity could provide enough SNR ($>10$) for timing and synchronisation.
@@
Dark count. The DCR of the detector used here is 500 Hz, and using SPCMs with DCR $<$ 100 Hz will result in a 7 dB increase in SNR. The reduced DCR will further decrease the QBER and increase the secure key rate.
Equivalent finite key length in real scenario. LEO’s intermittent communication window allows a limited length of key to be generated per satellite pass, which exacerbates the generation of secure keys as a significant portion of received signal must be sacrificed for accurate parameter estimation while retaining enough raw key for there to be a finite secure key after postprocessing. Considering the $QBER_{I}$ = 0.5%, the finite key length is about 6.5 Mbits with the elevation angle larger than 10 degrees. [35]. The aimed $R_{Q}$ is 400 MHz which could get at least a 12 dB SKL improvement over this 25 MHz system (assuming comparable $QBER$). Improving the quantum efficiency of the detector is another way to improve SKL but wouldn’t be too much.
Table 3: The parameter improvement in mission target
Description Parameter This Work Mission Target Gain
Quantum Source Repetition $R_{Q}$ 25 MHz 400 MHz $\uparrow$ 12 dB
Intrinsic QBER $QBER_{i}$ 1.5% 0.5% $\downarrow$ $\uparrow$
Dark Count (4 SPCMs) $DCR$ 2.3 kbps 2.3 kbps 0 dB [^7] This parameter should have improvement in the final system by replacing better component, but we leave them as the potential loss tolerance margin.
System Loss (785 nm) $L_{sys}$ 7.4 dB 3.8 dB $\downarrow$ 3.6 dB
Detector efficiency (785 nm) $L_{d}$ 60% (2.2 dB loss) 60% (2.2 dB loss) 0 dB77footnotemark: 7
Mean Photon Number $\mu$ 0.1 0.3744 $\uparrow$ 5.7 dB
In the comparison between this work and the mission target, there is a gain of 21.3 dB that can be achieved using the higher source repetition rate, higher $\mu$ (protected by decoy states) and lower system loss. Although we plan to use lower dark count detectors we also expect an increase of background light when outside the laboratory thus we expect broadly similar dark plus backround of around 2.3 kcps. Also the intrinsic QBER is expected to be slightly better but we will not consider it here as it’s non-linear effect on the SKR. With these assumptions we can project the result from this work to estimate the equivalent SKR in the real mission, as shown in Fig. 6. Using the gains in $\mu$ and reduced losses we can shift the maximum loss in Fig. 2 by 9.3 dB (without changing the shape of the curve) from 37.7 dB to 47 dB easily exceeding the maximum expected mission loss of 44.6 dB from table 1. The gain of 12 dB from increased repetition rate simple increases the key rate shifting the predicted key rate curve upwards. Although we only consider asymptotic key rates in this paper, the increase total key generated per satellite pass will increase our chances of also satisfying finite key requirements of security. As a comparison between the fiber link and free space link at the same distance of 1000 km, the SKR in satellite QKD is expected to be 3.47 x $10^{-4}$ per pulse, while in the Twin-Field QKD the SKR is 9.53 x $10^{-12}$ due to the extremely high loss of 156.5 dB [7].
Figure 6: Experimental result of SKR and the Equivalent SKR in the target mission. The black points shows the SKR evaluated in the experiment with a blue fit curve, and the orange dash line shows the estimated equivalent SKR when converting the experiment loss budget to the targeted mission budget. This involves shifting the curve horizontally by 9.3 dB taking into account reductions in system loss and increased $\mu$ and moving vertically by 12 dB to account for increases in repetition rate.
@@ Appendix A Channel Loss Estimation
Figure 7: The figure shows the distance and elevation angle change during the visibility window. The circle line shows the Elevation angle while the blue solid line shows the distance. The result is modeling based on the 3D Doppler model for the LEO [22]. The result is calculated with the value in the table 3.
Here, the channel loss modeling is given based on the diffraction limit as shown in Fig. 7. According to the analysis of the visibility window and the link efficiency, the elevation angle is configured larger than 10 degrees. The corresponding minimal and maximal distance is 500 km and 1700 km. Based on the profile of the Gaussian beam, the diffraction limited half angle with the threshold of $1/e^{2}$ is given by